# Denoising as Adaptation: Noise-Space Domain Adaptation for Image Restoration

### v1

@@ 5 Conclusion
In this work, we have presented a novel approach that harnesses the diffusion model as a proxy network to address the domain adaptation issues in image restoration tasks. Different from previous feature-space and pixel-space domain adaptation approaches, the proposed method adapts the restored results to their shared clean distribution in the pixel-wise noise space, resulting in significant low-level appearance improvements within a compact and stable training framework. To mitigate the shortcut issue arising from the joint training of the restoration and diffusion models, we randomly shuffle the channel index of two conditions and propose a residual-swapping contrastive learning strategy to prevent the diffusion model from discriminating the conditions based on the paired similarity. Furthermore, the proposed method can be extended by relaxing the input constraint of the diffusion model, introducing diverse unpaired clean images as denoising input. Experimental results have demonstrated the effectiveness of the proposed noise-space approach beyond existing feature-space and pixel-space methods on image restoration tasks. In the future, we plan to further investigate adapting the synthetic source domain to the real-world target domain using diffusion models, particularly in other dense prediction vision tasks.
@@
Appendix A More Implementation Details
A.1 Condition Evaluation on Diffusion Model
This work is inspired by the beneficial effects that favorable conditions facilitate the denoising process of the diffusion model, as shown in Fig 1(a). In this preliminary experiment, we first condition and train the diffusion model with an additional input in addition to its conventional input. Then, we test the noise prediction performance of this model under different qualities of the condition. To be specific, we corrupt the condition by adding the additive white Gaussian noise (AWGN) of noise level $\sigma\in[0,80]$ to its original clean images, which are performed on 1,000 images in the MS-COCO test dataset [79]. The noise prediction error of the diffusion model is evaluated using the mean square error (MSE) metric.
A.2 Comparison Settings
In comparison experiments, we mainly compare the proposed approach with three types of previous methods: domain adaptation methods, including DANN [19], DSN [22], PixelDA [27], and CyCADA [51]; unsupervised image restoration methods, including Ne2Ne [14], MaskedD [15], NLCL [16], SelfDeblur [5], and VDIP [17]; some representative supervised methods which serve as strong baselines in image restoration such as Restormer [31], to comprehensively evaluate generalization performance of different methods.
A.3 Scalability Evaluation
To provide a comprehensive evaluation of the proposed method, we apply six variants of the image restoration network in our experiments, including three variants of convolution-based network [82]: Unet-T (Tiny), Unet-S (Small), and Unet-B (Base); and three variants of Transformer-based network [32]: Uformer-T (Tiny), Uformer-S (Small), and Uformer-B (Base). These variants differ in the number of feature channels (C) and the count of layers at each encoder and decoder stage. The specific configurations, computational cost, and the parameter numbers are detailed below:
@@ A.3 Scalability Evaluation
Figure 9: Overview of different domain adaptation (DA) approaches. (a) Feature-space DA aligns the intermediate features across source and target domains. (b) Pixel-space DA translates source data to the “style" of the target domain through adversarial learning. (c) The proposed noise-space DA is specifically designed for image restoration. It gradually adapts the results from both source and target domains to the target clean image distribution, via multi-step denoising. Particularly, the function network represents a restorer in the context of image restoration.
Appendix B Discussion on Different Domain Adaptation Methods
@@
As discussed in Sec. 3.4, we described the effectiveness of the proposed method beyond the previous feature-space and pixel-space domain adaptation methods. We further show their specific framework in Fig. 9. In contrast to previous adaptation methods, our method is free to a domain classifier by introducing a meaningful diffusion loss function.
Appendix C More Visual Comparison Results
@@
We visualize more comparison results on the image denoising task in Fig. 10, image deraining task in Fig. 11, and image deblurring task in Fig. 12. We name the proposed method and its extension as ‘Ours’ and ‘Ours-Ex’, respectively.
Appendix D More Visual Results on Other Real-World Datasets
@@
To show the generalization ability of the proposed method, we also visualize the restored results of the proposed method on other real-world datasets [83, 45] in Fig. 13, Fig. 14, Fig. 15. These datasets were not encountered during the network’s training and fall outside the distribution of the trained datasets.
@@
Figure 10: Visual comparison of the image denoising task on SIDD test dataset [76].
@@
Figure 11: Visual comparison of the image deraining task on SPA test dataset [77].
@@
Figure 12: Visual comparison of the image deblurring task on RealBlur-J [78] test dataset.
@@
Figure 13: Visual results of the proposed method on DND real-world denoising test dataset [83].
@@
Figure 14: Visual results of the proposed method on DND real-world denoising test dataset [83].
@@
Figure 15: Visual results of the proposed method on ‘Real-Internet’ real-world deraining test dataset [45].

### v2

@@ 5 Conclusion
In this work, we have presented a novel approach that harnesses the diffusion model as a proxy network to address the domain adaptation issues in image restoration tasks. Different from previous feature-space and pixel-space adaptation approaches, the proposed method adapts the restored results to the target clean distribution in the pixel-wise noise space, resulting in significant low-level appearance improvements within a compact and stable training framework. To mitigate the shortcut issue arising from the joint training of the restoration and diffusion models, we randomly shuffle the channel index of two conditions and propose a residual-swapping contrastive learning strategy to prevent the model from discriminating the conditions based on the paired similarity. Furthermore, the proposed method can be extended by relaxing the input constraint of the diffusion model, introducing diverse unpaired clean images as denoising input. Experimental results demonstrate the effectiveness of our approach over feature-space and pixel-space domain adaptation methods, as well as its scalability surpassing that of self-supervised methods across a range of image restoration tasks.
@@
Appendix A1 Implementation Details
A1.1 Condition Evaluation on Diffusion Model
This work is inspired by the beneficial effects that favorable conditions facilitate the denoising process of the diffusion model, as shown in Fig 1(a). In this preliminary experiment, we first condition and train the diffusion model with an additional input in addition to its conventional input. Then, we test the noise prediction performance of this model under different conditions. To be specific, we corrupt the condition by adding the additive white Gaussian noise (AWGN) of noise level $\sigma\in[0,80]$ to its original clean images, which are performed on 1,000 images in the MS-COCO test dataset (Lin et al., 2014). The noise prediction error of the diffusion model is evaluated using the mean square error (MSE) metric.
A1.2 Comparison Settings
In comparison experiments, we mainly compare the proposed approach with three types of previous methods: domain adaptation methods, including DANN (Ganin & Lempitsky, 2015), DSN (Bousmalis et al., 2016), PixelDA (Bousmalis et al., 2017), and CyCADA (Hoffman et al., 2018); unsupervised image restoration methods, including Ne2Ne (Huang et al., 2021), MaskedD (Chen et al., 2023), NLCL (Ye et al., 2022), SelfDeblur (Ren et al., 2020), and VDIP (Huo et al., 2023); some representative supervised methods which serve as strong baselines in image restoration such as Restormer (Zamir et al., 2022), to comprehensively evaluate generalization performance of different methods. MaskedD (Chen et al., 2023) proposes masked training to enhance the generalization performance of denoising networks, showing the potential to be directly applicable to real-world scenarios. It shares the same goal with our work.
A1.3 Scalability Evaluation
To provide a comprehensive evaluation of the proposed method, we apply six variants of the image restoration network in our experiments, including three variants of convolution-based network (Ronneberger et al., 2015): Unet-T (Tiny), Unet-S (Small), and Unet-B (Base); and three variants of Transformer-based network (Wang et al., 2022): Uformer-T (Tiny), Uformer-S (Small), and Uformer-B (Base). These variants differ in the number of feature channels (C) and the count of layers at each encoder and decoder stage. The specific configurations, computational cost, and the parameter numbers are detailed below:
@@
Figure A1: Overview of different domain adaptation (DA) approaches. (a) Feature-space DA aligns the intermediate features across source and target domains. (b) Pixel-space DA translates source data to the “style" of the target domain through adversarial learning. (c) The proposed noise-space DA is specifically designed for image restoration. It gradually adapts the results from both source and target domains to the target clean image distribution, via multi-step denoising. Particularly, the function network represents a restoration network in the context of image restoration.
Appendix A2 Discussion on Different Domain Adaptation Methods
@@
As discussed in Sec. 3.4, we described the effectiveness of the proposed method beyond the previous feature-space and pixel-space domain adaptation methods. We further show their specific framework in Fig. A1. In contrast to previous adaptation methods, our method is free to a domain classifier or discriminator by introducing a meaningful diffusion loss function.
Appendix A3 Additional Analysis of the Ablations
We provided an ablation study to show the necessity of real data, in which only the synthetic or real data conditions onto the diffusion model. The quantitative results of the SIDD test dataset are listed in Table. 4. It is noteworthy that both synthetic and real data are essential for effective domain adaptation in diffusion models. Omitting either type results in a significant decline in real-world performance. In particular, when real data is excluded, the performance nearly degrades to the level of a Vanilla model. We further analyze the necessity of each condition as follows: (1) Real data typically acts as a “bad” condition that introduces extra noises to the diffusion model, because the restoration network cannot restore it well under the domain gap. Consequently, valid and strong diffusion loss would backpropagate to the restoration network, promoting it learns to provide “good” conditions. As a benefit of the proposed strategies to eliminate the shortcut, the model progressively adapts the real data into the target clean distribution in a multi-step denoising manner. (2) Synthetic data in two conditions can provide useful guidance in the early training stage, ensuring the diffusion model continuously focuses on these condition channels.
Appendix A4 Additional Comparison Results
Table A1: Quantitative metrics of the proposed method (Ours) and its extension on unpaired condition case (Our-Ex). The results are formed with PSNR/SSIM/LPIPS. The best and second best scores are highlighted and underlined.
. Task Ours Ours-Ex Denoising 34.71/0.9202/0.0903 33.44/0.8938/0.1064 Deraining 34.39/0.9571/0.0462 34.20/0.9587/0.0444 Deblurring 26.46/0.8048/0.1363 26.44/0.8030/0.1313
A4.1 Extension
As mentioned in Sec. 3.2, our method can extend to the unpaired condition case by relaxing the diffusion’s input with the image from other clean datasets. Thus, the shortcut issue can be potentially eliminated since the trivial solutions such as matching the pixel’s similarity between input and condition do not exist. Such an extension keeps the channel shuffling layer but is free to the residual swapping contrastive learning. We show the quantitative evaluation in Tab. A1. The results demonstrate that although the condition and diffusion input are unpaired, our method can still learn to adapt the restored results from the synthetic and real-world domains to the clean image distribution, which also complements the restoration performance of the paired solution in some tasks like deraining and deblurring.
A4.2 More Advanced Restoration Networks
As discussed in Sec. 4.3, the proposed domain adaptation method offers strong scalability across various image restoration networks. Additionally, by employing more advanced restoration networks with the proposed denoising as adaptation (Ours*), the performance can be further improved, yielding results that are more perceptually aligned with the ground truth as illustrated in Fig. A2. The complexity comparison of different image restoration networks is listed in Table A2.
Table A2: Complexity comparison of the image restoration methods: Parameter (M), GMACs (G).
Metrics C2N AP-BSN Restormer Selfdeblur MaskedD VDIP Ours Ours*
Parameter 164.57 3.10 26.09 3.10 12.00 3.00 8.56 65.19
GMACs 280.48 60.62 35.24 23.83 188.03 23.77 12.48 22.46
Figure A2: Visual results on detailed textures and high-frequency components. The proposed method serves as a general learning strategy for the image restoration task, offering scalability across different restoration networks. It also enables performance improvements as the complexity of the restoration network increases (Ours*).
A4.3 Additional Visual Comparison Results
@@
We visualize more comparison results on the image denoising task in Fig. A3, image deraining task in Fig. A4, and image deblurring task in Fig. A5. In particular, we name the proposed method and its extension as ‘Ours’ and ‘Ours-Ex’, respectively.
A4.4 Additional Visual Results on Other Real-World Datasets
@@
To show the generalization ability of the proposed method, we also visualize the restored results of the proposed method on other real-world datasets (Plotz & Roth, 2017; Yang et al., 2017) in Fig. A6, Fig. A7, Fig. A8. Please note that these datasets were not seen during the network’s training and fall outside the distribution of the trained datasets.
@@
Figure A3: Visual comparison of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018).
@@
Figure A4: Visual comparison of the image deraining task on SPA test dataset (Wang et al., 2019).
@@
Figure A5: Visual comparison of the image deblurring task on RealBlur-J (Rim et al., 2020) test dataset.
@@
Figure A6: Visual results of the proposed method on DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A7: Visual results of the proposed method on DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A8: Visual results of the proposed method on ‘Real-Internet’ real-world deraining test dataset (Yang et al., 2017).